# NeuTSFlow: Modeling Continuous Functions Behind Time Series Forecasting

### v1

@@ 4 Experiment
4.1 Experiment Settings
To demonstrate the superiority of our model, we compared it against several representative baselines, including Transformer-based models such as iTransformer [19], FedFormer [32], Pathformer [4], and PatchTST [22]. Additionally, we evaluated our model against MLP-based approaches like DLinear [30] and TimesMixer [27], as well as other established models such as Autoformer [3], Pyraformer [15], and Informer [31]. Due to space limitations and the need for prioritization, we selected the six most representative models for MSE comparison. Our experiments leverage eight real-world datasets, including ECL, four subsets of ETT, Exchange, Traffic, and Weather, which were previously employed in Autoformer [3]. The complete results are provided in Appendix A.4. A detailed description of the relevant models and datasets can be found in Appendix A.5.
@@ 4.2 Time Series Super-resolution
Figure 3: From left to right are the super-resolution results of NeuTSFlow, iTransformer, and DLinear
@@
Table 1: Time Series Super resolution
@@
Task Setup To construct a realistic dataset pair with high-sampling-rate (HSR) and low-sampling-rate (LSR) characteristics for time series analysis, we employ established benchmark datasets—such as the ETTh1 dataset—as the HSR reference. The corresponding LSR dataset is generated through a fixed-interval decimation process, where data points are subsampled at regular temporal intervals. Specifically, the LSR input sequence length is set to 24 time steps, derived from an HSR input sequence length of 96 time steps via a temporal decimation factor of 4. Both HSR and LSR sequences are strictly aligned to share identical temporal start and end, ensuring phase consistency and direct comparability across resolutions. This alignment is critical to preserve the temporal correspondence between datasets, avoiding misalignment errors that could compromise model training.
Table 2: Cross-Resolution Temporal Learning
@@ 4.4 Conventional Forecasting Task
Task Results As shown in Table LABEL:forecasting, NeuTSFlow demonstrates significant overall advantages in traditional time series forecasting tasks. It performs exceptionally well on datasets with strong periodicity and structured patterns (e.g., the ETT series, Exchange, and ECL). While NeuTSFlow does not rank first on some other datasets (placing second in both cases), it remains highly competitive across the board.
Table 3: Conventional Forecasting Task
@@ 4.5 Ablation studies
To validate the contribution of key modules in NeuTSFlow, we conduct ablation experiments by removing four critical components: flow matching, neural operator, normalization layer, and spectral decomposition layer. The experiments are performed on three representative datasets across three tasks. The results are reported in Table LABEL:ablation as MSE values. "W/O neural operator" indicates replacing the Marginal Velocity Field Model with the DLinear model. "W/O flow matching" refers to using only the Marginal Velocity Field Model for prediction.
As shown in Table LABEL:ablation, the removal of the Neural Operator results in the largest performance drop, underscoring its critical role in modeling continuous temporal dynamics and the importance of a function-based perspective. Similarly, the removal of Flow Matching leads to the second-largest impact, highlighting its effectiveness in capturing transitions between function families. The removal of the Normalization Layer primarily affects non-stationary data, demonstrating its significance in managing dynamic variability. Overall, the Neural Operator and Flow Matching are the core innovations behind NeuTSFlow’s superior performance. These results validate the utility of each component and support the scientific rationale for integrating discrete-domain operations with continuous-time modeling, establishing a robust framework for time seires forecasting.
Table 4: Ablation Study Results on Key Modules
@@
Cross-Resolution Temporal Learning ETTh1 0.370 0.442 0.427 0.405 0.418
@@
Time Series Super-Resolution ETTh1 0.0842 0.1182 0.1021 0.0997 0.0910

### v2

@@ 4 Experiment
We empirically validate NeuTSFlow on multiple benchmark datasets and forecasting tasks. The goal is to assess its performance in modeling function-level transitions, as well as its ability to generalize across varying temporal resolutions and domains.
Experiment Settings. To demonstrate the superiority of our model, we compared it against several representative baselines, including Transformer-based models such as iTransformer [19], FedFormer [32], Pathformer [4], and PatchTST [22]. Additionally, we evaluated our model against MLP-based approaches like DLinear [30] and TimesMixer [27], as well as other established models such as Autoformer [3], Pyraformer [15], and Informer [31]. Due to space limitations and the need for prioritization, we selected the six most representative models for MSE comparison. Our experiments use eight real-world datasets, including ECL, four subsets of ETT, Exchange, Traffic, and Weather, which were previously employed in Autoformer [3]. The full results are provided in Appendix A.4. A detailed description of the relevant models and datasets can be found in Appendix A.5.
@@ 4.1 Time Series Super-resolution
Task Setup To construct a realistic dataset pair with high-sampling-rate (HSR) and low-sampling-rate (LSR) characteristics for time series analysis, we employ established benchmark datasets—such as the ETTh1 dataset—as the HSR reference. The corresponding LSR dataset is generated through a fixed-interval decimation process, where data points are subsampled at regular temporal intervals.
Figure 3: From left to right: super-resolution results produced by DLinear, iTransformer, and NeuTSFlow.
Specifically, the LSR input sequence length is set to 24 time steps, derived from an HSR input sequence length of 96 time steps via a temporal decimation factor of 4. Both HSR and LSR sequences are strictly aligned to share identical temporal start and end, ensuring phase consistency and direct comparability across resolutions. This alignment is critical to preserve the temporal correspondence between datasets, avoiding misalignment errors that could compromise model training.
Table 1: Time series super-resolution performance (MSE) across eight benchmark datasets. Lower values indicate better performance. NeuTSFlow consistently achieves the best results, ranking first on all datasets and outperforming state-of-the-art baselines by a significant margin.
@@
Table 2: Evaluation of model performance on the Cross-Resolution Temporal Learning (CRTL) task. NeuTSFlow has the best results across all datasets, indicating its effectiveness in learning resolution-agnostic functional representations.
@@ 4.3 Conventional Forecasting Task
Task Results As shown in Table LABEL:forecasting, NeuTSFlow outperforms in traditional time series forecasting tasks, especially on datasets with strong periodicity and structured patterns like ETT, Exchange, and ECL. While it does not always rank first, it remains highly competitive across all datasets, showcasing its robustness and versatility.
Table 3: Mean Squared Error (MSE) results for the conventional time series forecasting task across eight benchmark datasets. NeuTSFlow consistently ranks among the top-performing models, achieving the lowest average rank and outperforming established Transformer-based and MLP-based baselines on most datasets.
@@ 4.4 Ablation studies
To validate the contribution of key modules in NeuTSFlow, we conduct ablation experiments by removing four critical components: flow matching, neural operator, normalization layer, and spectral decomposition layer. The experiments are performed on three representative datasets across three tasks. The results are reported in Table LABEL:ablation as MSE values. "W/O neural operator" indicates replacing the Marginal Velocity Field Model with the DLinear model. "W/O flow matching" refers to using only the Marginal Velocity Field Model for prediction. As shown in Table LABEL:ablation, the removal of the Neural Operator results in the largest performance drop, underscoring its critical role in modeling continuous temporal dynamics and the importance of a function-based perspective. Similarly, the removal of Flow Matching leads to the second-largest impact, highlighting its effectiveness in capturing transitions between function families. The removal of the Normalization Layer primarily affects non-stationary data, demonstrating its significance in managing dynamic variability. Overall, the Neural Operator and Flow Matching are the core innovations behind NeuTSFlow’s superior performance. These results validate the utility of each component and support the rationale for integrating discrete-domain operations with continuous-time modeling, yielding to a robust framework for time series forecasting.
Table 4: Ablation results across three tasks. Neural operators and flow matching are most critical, while normalization and spectral decomposition also boost performance. CFT=Conventional Forecasting Task, CRTL= Cross-Resolution Temporal Learning, and TSSR=Time Series Super-Resolution
@@
CRTL ETTh1 0.370 0.442 0.427 0.405 0.418
@@
TSSR ETTh1 0.0842 0.1182 0.1021 0.0997 0.0910
@@ 5 Conclusion
6 Acknowledgement
CWC is supported by the Swiss National Science Foundation under grant number 20HW-1 220785. AIAR gratefully acknowledges the support of the Yau Mathematical Sciences Center, Tsinghua University. This work is also supported by the Tsinghua University Dushi Program.